ZombieLover:

Add the following polynomials (Answer is equation) Find the sum of 7a^3 + 14a + 12 and -6a^3 + 12a^2 -7.

boredfr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @boredfr @Justus \(\color{#0cbb34}{\text{End of Quote}}\) @darkknight

darkknight:

Find the sum of 7a^3 + 14a + 12 and -6a^3 + 12a^2 -7. combine like terms, 7a^3 - 6a^3= a^3 14a + 0a = 14a +12a^2 + 0a^2 = 12a^2 12 + (-7) = 5 a^3 +12a^2+14a+5